A principal of $5000 was invested in a savings account for 4 years. If the interest earned for the period was $400, what was the annual interest rate?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

A principal of $5000 was invested in a savings account for 4 years. If the interest earned for the period was $400, what was the annual interest rate?
-----------------
I = Prt
400 = 5000*r*4
r = 400/(4*5000)
r = 100/5000
r = 1/50
r = 0.02 or 2%
=================
Cheers,
Stan H.

Answer by checkley71(8403)

i=prt where i=interest, p=principle, r=interest rate & t=number of years.
400=5000*r*4
400=20000r
r=400/20000
r=.02 or 2% is the interest rate. This assumes the rate is for the 4 years NOT an annual interest rate.
An annual interest rate formula is:
i=p(1+r)^t-p
400=5000(1+r)^4-5000
400+5000=5000(1+r)^4
5400=5000(1+r)^4
5400/5000=(1+r)^4
1.08=(1+r)^4
4th root of 1.08=1+r
1.019426547=1+r
r=1.0194-1
r=.0194 or 1.94% if compounded annually.